47 What is the conce to air which protects the bulk metal from Aluminum forms a layer of aluminum oxide when exposed further corrosion , calculate AS° for this reaction. Using the thermodynamic data provided below S°/K-mol) 28.3 205.0 z0g) 20(s) 50.99 a. 182.3 J/K mol b. 131.5 J/K mol c -182.3 J/K mol d. -626.2 J/K.mol e -802.9 J/K-mol at is Kyp for this compound? e. 2.1 x 106 41. The molar solubility of tin(II) iodide is 1.28 × 10-2 mol/L. wh 4.2×10-6 a. 8.4 x 10 b, 1.28 × 10-2 e. d. 1.6×10-4 42, Nitrosyl chloride (NOCl) decomposes at elevated temperatures according to the equation 2NOCI(g) → 2NO(g) + Cl2(g). Calculate Kp for this reaction at 227°C ( 110 kJ/mol, AS 128 J/K mol) 1.59 × 10-2 c. 62.8 e. 3.20 x 10 b. 2.10 x 10-7 d. 4.90 x 106 Help as soon as possible
Answer #1

40. D

for 4Al(s) + 3O2(g) ---> 2Al2O3(s)

We can calculate delta S° by the equation; delta S° = S°products - S°reactants.

Therefore delta S° = 2 x 50.99 – [4 x 28.3 + 3 x 205.0] = 101.98 – [113.2 + 615] = 101.98 – 728.2 = -626.2 J/Kmol